From GlueXWiki
Jump to: navigation, search

October 31, 2013 FDC meeting


  1. Installation [1]
  2. Engineering (Bill)
  3. Electronics (Chris, Nick) list of cablesNick's presentation about cable locations on platform
  4. Tests in EEL126 FDC E-log (Beni)
    • New F1TDC
  5. Other


Participants: Fernando, Bill, Eugene, Glenn, Tom, Tim, Chris, Simon, Nick, Vlad, Beni, and Lubomir. Long meeting discussing the survey/alignments most of the time.


- The results from the survey of the FDC on the floor in the Hall are attached (still preliminary). The main conclusions: flatness of all the packages is OK (max deviation from the best fitted plane is 0.022"=0.56 mm). The centering of all the packages is perfect. The main concern is that the tilt (pitch and yaw) of the packages changed if compared to the Blue Crab measurements, which means FDC is flexible to this extend. Mostly the first and the last packages had deviations up to 0.66 mm. We just need to know the tilts, but the question is what will happen when FDC is moved inside the magnet. Eugene's point: it's very difficult to do tilt corrections using the beam data, we need to know the z positions better than 0.5mm and most importantly be sure it doesn't change when inserting the FDC. Independently, Bill and Lubomir measured the distance between the rails on the FDC and mesh carts and at the upstream end of BCAL. The rails in BCAL and the mesh cart matched very well, but the distance b/n rails on the FDC cart was by 0.8 mm different at the two ends. Tim and Tom say this doesn't cause additional friction and therefore deformation (i.e. it doesn't explain the above discrepancy b/n Blue Crab and Hall survey).

- Therefore Tim, Bill, Tom proposed (and everybody agreed) to try to do survey when the FDC is in its final position connected with cables. Most of the discussions: if it is possible, and how to do it. For the outer packages everybody agrees it is doable. For the inner packages there must be different solutions, but for now Bill will start looking for non-magnetic targets to be installed in points visible from the sides when the FDC is inside the magnet.

- The short term plan: since we think the survey of the FDC in final position is doable we can go ahead and put FDC back on the platform (on Friday). Then we will move it in and out of the magnet and do the survey in both positions. Tom will have people working on the cable mesh. After it's done we can start cabling.

- FDC was moved to the Hall a week ago (Oct 24) (see pictures attached). Since then it is continuously flashed with gas and it bubbles normally. We tried to insert FDC in the magnet and it was almost flash with the upstream end of BCAL but got stuck there. The problem was a HV connector caught the head of a screw on the rails (last picture). The size of the heads turned out to be bigger for the rails inside the BCAL. These screws will be replaced (one by one); at the same time the the plates holding the HV connector were modified the move the connector close to the beam. Since Monday the FDC is down on the floor for the survey. It was not possible to do it on the platform. There Casey also connected the cooling manifolds with tubes.


- Apart from the above discussions Eugene asked about the status with the chilled water and the cooling system. Bill: the piping is done and they are now testing it. Eugene: what is the water pressure (100psi) and what will happen if the piping breaks (very bad).


- Nick made a presentation (linked above) about where the cables are on the platform and how the boxes are labeled. So far the signal cables for the first package and the LV cables are sorted in this way. We will first connect one quadrant of package #1, then same quadrant for package #2 ... starting with the LV cables. Therefore we will need the signal cables for all the packages almost from the beginning, but this will be done soon (Nick).

- About 10 cables from each package are missing, Tom will check if those are in the ESB used there for the mock-up.

Tests at EEL126

- Beni was able to use the pulser on the new F1TDC: the width of the timing distributions with the pulser is ~0.3 ns.